1.Find the potential at a distance of 0.07 m from a -6 nC charge.


2.What would be the PE of a -5 nC charge placed at this point?


3.What is the electric field strength at a point 40 cm from the a charge of 5x10^-9 C?

1
Expert's answer
2022-03-09T12:30:44-0500

(a) potential

"V=\frac{kq}{r}"


"V=-\frac{9\times10^9\times6\times10^{-9}}{0.07}=-771.42volt"

(b) PE

"U=-qV"


"U=((-5\times10^{-9})\times(-771.42))=3.85\times10^{-6}J"

(C)

Electric field

"E=\frac{kq}{r^2}"

"E=\frac{9\times10^9\times5\times10^{-9}}{0.4^2}=281.25N/C"